# Non-toric brane webs, Calabi–Yau 3-folds, and 5d SCFTs

### v1

@@ 3.1.2. Log and open Calabi–Yau surfaces
If $(Y,D)$ is a log Calabi-Yau surface and $p$ is a singular point of $D$, then, denoting $\widetilde{Y}$ the blow-up of $Y$ at $p$ and $\widetilde{D}$ the strict transform of $D$ in $\widetilde{Y}$, the pair $(\widetilde{Y},\widetilde{D})$ is a log Calabi–Yau surface. We say that $(\widetilde{Y},\widetilde{D})$ is obtained from $(Y,D)$ by corner blow-up. The open Calabi–Yau surface $U=Y\setminus D$ does not change under a corner blow-up.
@@
If $(Y,D)$ is a log Calabi-Yau surface and $p$ is a smooth point of $D$, then, denoting $\widetilde{Y}$ the blow-up of $Y$ at $p$ and $\widetilde{D}$ the strict transform of $D$ in $\widetilde{Y}$, the pair $(\widetilde{Y},\widetilde{D})$ is a log Calabi–Yau surface. We say that $(\widetilde{Y},\widetilde{D})$ is obtained from $(Y,D)$ by interior blow-up. The open Calabi–Yau surface $U=Y\setminus D$ changes under an interior blow-up. In particular, interior blow-ups of a toric log Calabi–Yau surface are examples of non-toric log Calabi–Yau surfaces. According to [47, Proposition 1.3], for every log Calabi–Yau surface $(Y,D)$, there exists a log Calabi–Yau surface $(\widetilde{Y},\widetilde{D})$ obtained from $(Y,D)$ by corner blow-ups, and such that $(\widetilde{Y},\widetilde{D})$ admits a toric model, that is, a morphism $(\widetilde{Y},\widetilde{D})\rightarrow(\overline{Y},\overline{D})$ that is a composition of interior blow-ups to a toric surface $(\overline{Y},\overline{D})$.
@@ 3.1.3. Asymptotic configurations of $7$ branes and open Calabi–Yau surfaces
[…]
Figure 3.2. Elementary cluster transformation: after a blow-up producing the exceptional curve $E_{i}^{+}$, contract the $(-1)$-curve $E_{i}^{-}$.
Without loss of generality, we can assume that the fan of $\overline{Y}_{A}$ contains the line $\mathbb{R}(p_{i},q_{i})$. Then, the projection $\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}/\mathbb{R}(p_{i},q_{i})=\mathbb{R}$ induces a toric morphism $\nu:\overline{Y}_{A}\rightarrow\mathbb{P}^{1}$, which is generically a $\mathbb{P}^{1}$-fibration. Moreover, the irreducible components $D_{i}^{+}$ and $D_{i}^{-}$ of $\overline{D}$ corresponding respectively to the rays $\mathbb{R}_{\geq 0}(p_{i},q_{i})$ and $\mathbb{R}_{\geq 0}(-p_{i},-q_{i})$ of the fan, are sections of $\nu$, see Figure 3.2. By Construction 3.10, $Y_{A}$ is obtained from the toric surface $\overline{Y}_{A}$ by blowing up points $y_{j}$ on the toric boundary for each $1\leq j\leq Q$. In particular, blowing up the point $y_{i}\in D_{i}^{+}$ creates an exceptional $(-1)$-curve $E_{i}^{+}$ in $Y_{A}$. The strict transform of the $\mathbb{P}^{1}$-fiber of $\nu$ passing through $x_{i}$ is also a $(-1)$-curve $E_{i}^{-}$. Hence, one can construct a new toric surface $Y_{A_{i}^{\prime}}$ from $Y_{A}$ by contracting all the exceptional curves coming from blowing up the points $y_{j}$ with $j\neq i$, and the $(-1)$-curve $E_{i}^{-}$. The induced birational map $\overline{Y}_{A}\dashrightarrow\overline{Y}_{A_{i}}$ is an example of “elementary cluster transformation” as in [46, §3.1], and it follows from [46] that the fan of the toric surface $\overline{Y}_{A_{i}}$ is obtained from the fan of $\overline{Y}_{A}$ by the piecewise-linear transformation given by (3.1). In particular, the toric surface $Y_{A_{i}^{\prime}}$ can be taken as the starting point of the Construction 3.10 for the asymptotic configurations $A_{i}^{\prime}$ obtained from $A$ by Hanany–Witten move of the $i$-th 7-brane. The corresponding log Calabi–Yau surface $(Y_{A_{i}^{\prime}},D_{A_{i}^{\prime}})$ coincides with $(Y_{A},D_{A})$, and so the corresponding open Calabi–Yau surface $U_{A_{i}^{\prime}}$ coincides with $U_{A}$. Moreover, the elementary cluster transformation restricted to the tori $(\mathbb{C}^{\star})^{2}_{A}$ and $(\mathbb{C}^{\star})^{2}_{A_{i}^{\prime}}$ defines a birational transformation
@@
for all $1\leq i,j\leq Q$. It follows from (3.1) that Hanany–Witten moves for asymptotic configurations of 7-branes are exactly mutations of seeds in the theory of cluster algebras [40]. The seed $\mathbf{s}$ defines a $Q$-dimensional $\mathcal{X}$ cluster variety with a natural Poisson structure induced by $\omega$ [39] and a cluster chart $(\mathbb{C}^{\star})^{Q}\subset\mathcal{X}$. A seed mutation does not change the $\mathcal{X}$ cluster variety but changes the cluster chart by an elementary cluster modification. The skew-symmetric form $\omega$ has rank $2$, and it follows that the symplectic leaves of the Poisson structure on the $\mathcal{X}$ cluster variety are 2-dimensional. By [46], the symplectic leaves of the $\mathcal{X}$ cluster variety coincide up to codimension $2$ with the open Calabi–Yau surfaces deformation-equivalent to the open Calabi–Yau surface $U$ associated to the asymptotic configuration of 7-branes by Construction 3.10.
@@
To prove that the map $F$ is injective, let $A_{1}$ and $A_{2}$ be two asymptotic configurations of 7-branes such that the corresponding open Calabi–Yau surfaces $U_{A_{1}}$ and $U_{A_{2}}$ are isomorphic. For every $k\in\{1,2\}$, we have $U_{A_{k}}=Y_{A_{k}}\setminus D_{A_{k}}$, where $(Y_{A_{k}},D_{A_{k}})$ is a log Calabi–Yau surface defined via a toric model $(Y_{A_{k}},D_{A_{k}})\rightarrow(\overline{Y}_{A_{k}},\overline{D}_{A_{k}})$. By [55, Proposition 3.27] (see also [21, Theorem 1]), any two toric models of a log Calabi–Yau surface are connected by a sequence of corner blow-ups/blow-downs and elementary cluster transformations. Since elementary cluster transformations correspond to Hanany–Witten moves at the level of the asymptotic configurations of 7-branes, it follows that $A_{1}$ and $A_{2}$ are equivalent in the sense of Definition 3.2. This concludes the proof that the map $F$ is injective. ∎
@@ 3.2. Webs of 5-branes with 7-branes
[…]
Figure 3.3. A web of 5-branes with 7-branes
@@
Let $\overline{W}\subset\mathbb{R}^{2}$ be a web of 5-branes, and $\mathbf{a}:=((a_{e,i})_{1\leq i\leq n_{e}})_{E\in L(\overline{W})}$ a 7-brane data for $\overline{W}$. We define an integral affine manifold with singularities $B$, by introducing for every leg $e$ of $\overline{W}$, $n_{e}$ focus-focus singularities $x_{e,1},\dots,x_{e,n_{e}}$ on $e$ with monodromy-invariant direction the direction $(p_{e},q_{e})$ of $e$, ordered as in Figure 3.3. Let $W\subset B$ be the embedded graph in $B$ obtained from $\overline{W}$ as follows:
@@
    Finally, we assign weight $w_{e}$ to the edge $[v,x_{e,1}]$, weights $w_{e}-\sum_{j=1}^{i}a_{e,j}$ to the edge $[x_{e,i},x_{e,i+1}]$ for all $11\leq i\leq n_{e}-$ 1, and weight $w_{e}-\sum_{j=1}^{n_{e}}a_{e,j}$ to the leg $x_{e,n_{e}}+\mathbb{R}_{\geq 0}(p_{e},q_{e})$ if $\sum_{i=1}^{n_{e}}a_{e,i}<w_{e}$.
@@ 3.3. Asymptotic webs of 5-branes with 7-branes and log Calabi–Yau surfaces
Recall from §2.2 that the asymptotic 5-brane $\overline{W}^{\mathrm{asym}}$ determines a toric surface $(\overline{Y},\overline{D})$ with an ample line bundle $\overline{L}$. Every leg $e\in L(\overline{W}^{\mathrm{asym}})$ corresponds to a toric irreducible component $\overline{D}_{e}$ of $\overline{D}$. Denote by $\overline{D}_{e}^{\circ}$ the open stratum of $\overline{D}_{e}$, that is, the complement in $\overline{D}_{e}$ of the 0-dimensional strata of $\overline{D}$. For every leg $e\in L(\overline{W}^{\mathrm{asym}})$, pick $n_{e}$ distinct points $y_{e,i}$ of $\overline{D}_{e}^{\circ}$, where $1\leq i\leq n_{e}$. Let $Y$ be the projective surface obtained by blowing up the toric surface $\overline{Y}$ at the points $y_{e,i}$ for all $e\in L(\overline{W}^{\mathrm{asym}})$ and $1\leq i\leq n_{e}$. Then, denoting by $D$ the strict transform of $\overline{D}$ in $Y$, $(Y,D)$ is a log Calabi–Yau surface. Finally, we define a line bundle $L$ on the surface $Y$ by
@@
where $\pi:Y\rightarrow\overline{Y}$ is the blow-up map, and $E_{e,i}=\pi^{-1}(y_{e,i})$ are the exceptional divisors in $Y$ over the points $y_{e,i}$ – see Figure 3.4.
[…]
Figure 3.4. The log Calabi–Yau surface associated to an asymptotic web of 5-branes with 7-branes.
@@
While for Type I webs $W^{\mathrm{asym}}$, we do not impose any restrictions on the configuration of the points $y_{e,i}$, for Type II webs we require that the points $y_{e,i}$ to satisfy the Menelaus condition with respect to $L$, that is, that the line bundle $L|_{D}$ is trivial – see for example [82] for a previous use of this terminology.
@@
Since $D$ is a nodal curve of arithmetic genus one, the group of $\mathrm{Pic}^{0}(D)$ of line bundles on $D$ with degree 00 on every irreducible component of $D$ is isomorphic to $\mathbb{C}^{\star}$. By Equation 3.5, we have $L_{D}\in\mathrm{Pic}^{0}(D)$, and so the Menelaus condition can be written as $L|_{D}=1\in\mathbb{C}^{\star}$. Recall that the points $y_{e,i}$ lie in the interior of the big torus orbits in the boundary components $\overline{D}_{e}$’s of $\overline{D}$. Denote by $z_{e}$ a coordinate on the component $\overline{D}_{e}$, with the convention that we choose an orientation $\overline{D}$ and denote the coordinates of the two nodal points intersecting adjacent components by 00 and $\infty$ respectively, following this orientation. Denoting by $z_{e}(y_{e,i})\in\mathbb{C}^{*}$ the coordinates of $y_{e,i}$, the Menelaus condition amounts to the requirement:
@@ 3.4. Webs of $5$-branes with $7$-branes and degenerations of log Calabi–Yau surfaces
In this section, after defining integral webs of $5$ branes with $7$ branes $W$, we explain how to construct a one-parameter degeneration $(\mathcal{Y},\mathcal{D})\to\mathbb{C}$ of the log Calabi–Yau surface $(Y,D)$, associated to $W^{\mathrm{asym}}$ as in §3.2. We also explain how to construct a degeneration $\mathcal{L}$ of the line bundle $L$ on $Y$.
Let $W$ be an integral web of $5$ branes with $7$ branes, obtained from a web of $5$ branes $\overline{W}$. As explained in §2.2.3, the web $\overline{W}$ defines a polarized toric degeneration
@@
with central fiber the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ defined by the associated asymptotic web $\overline{W}^{\mathrm{asym}}$ as in §2.2.2.
@@
where $p:\mathcal{Y}\rightarrow\overline{\mathcal{Y}}$ is the blow-up map, and $\mathcal{E}_{e,i}$ are the exceptional divisors in $\mathcal{Y}$ over the images of the sections $\sigma_{e,i}$. By construction, the restriction of $\mathcal{L}$ to a general fiber $(\mathcal{Y}_{t},\mathcal{D}_{t})\simeq(Y,D)$ is isomorphic to the line bundle $\overline{L}$ defined in Equation (3.3). From now on we refer to

### v2

@@ 3.1.2. Log and open Calabi–Yau surfaces
[…]
@@
[…]
@@ 3.1.3. Asymptotic configurations of $7$ branes and open Calabi–Yau surfaces
Figure 5. Elementary cluster transformation: after a blow-up producing the exceptional curve $E_{i}^{+}$, contract the (1)(-1)-curve $E_{i}^{-}$.
Without loss of generality, we can assume that the fan of $\overline{Y}_{A}$ contains the line $\mathbb{R}(p_{i},q_{i})$. Then, the projection $\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}/\mathbb{R}(p_{i},q_{i})=\mathbb{R}$ induces a toric morphism $\nu:\overline{Y}_{A}\rightarrow\mathbb{P}^{1}$, which is generically a $\mathbb{P}^{1}$-fibration. Moreover, the irreducible components $D_{i}^{+}$ and $D_{i}^{-}$ of $\overline{D}$ corresponding respectively to the rays $\mathbb{R}_{\geq 0}(p_{i},q_{i})$ and $\mathbb{R}_{\geq 0}(-p_{i},-q_{i})$ of the fan, are sections of $\nu$, see Figure 5. By Construction 3.10, $Y_{A}$ is obtained from the toric surface $\overline{Y}_{A}$ by blowing up points $y_{j}$ on the toric boundary for each $1\leq j\leq Q$. In particular, blowing up the point $y_{i}\in D_{i}^{+}$ creates an exceptional (1)(-1)-curve $E_{i}^{+}$ in $Y_{A}$. The strict transform of the $\mathbb{P}^{1}$-fiber of $\nu$ passing through $x_{i}$ is also a (1)(-1)-curve $E_{i}^{-}$. Hence, one can construct a new toric surface $Y_{A_{i}^{\prime}}$ from $Y_{A}$ by contracting all the exceptional curves coming from blowing up the points $y_{j}$ with $j\neq i$, and the (1)(-1)-curve $E_{i}^{-}$. The induced birational map $\overline{Y}_{A}\dashrightarrow\overline{Y}_{A_{i}}$ is an example of “elementary cluster transformation” as in [49, §3.1], and it follows from [49] that the fan of the toric surface $\overline{Y}_{A_{i}}$ is obtained from the fan of $\overline{Y}_{A}$ by the piecewise-linear transformation given by (3.1). In particular, the toric surface $Y_{A_{i}^{\prime}}$ can be taken as the starting point of the Construction 3.10 for the asymptotic configurations $A_{i}^{\prime}$ obtained from $A$ by Hanany–Witten move of the $i$-th 7-brane. The corresponding log Calabi–Yau surface $(Y_{A_{i}^{\prime}},D_{A_{i}^{\prime}})$ coincides with $(Y_{A},D_{A})$, and so the corresponding open Calabi–Yau surface $U_{A_{i}^{\prime}}$ coincides with $U_{A}$. Moreover, the elementary cluster transformation restricted to the tori $(\mathbb{C}^{\star})^{2}_{A}$ and $(\mathbb{C}^{\star})^{2}_{A_{i}^{\prime}}$ defines a birational transformation
@@
for all $1\leq i,j\leq Q$. It follows from (3.1) that Hanany–Witten moves for asymptotic configurations of 7-branes are exactly mutations of seeds in the theory of cluster algebras [43]. The seed $\mathbf{s}$ defines a $Q$-dimensional $\mathcal{X}$ cluster variety with a natural Poisson structure induced by $\omega$ [42] and a cluster chart $(\mathbb{C}^{\star})^{Q}\subset\mathcal{X}$. A seed mutation does not change the $\mathcal{X}$ cluster variety but changes the cluster chart by an elementary cluster modification. The skew-symmetric form $\omega$ has rank $2$, and so the symplectic leaves of the Poisson structure on the $\mathcal{X}$ cluster variety are 2-dimensional. By [49], the symplectic leaves of the $\mathcal{X}$ cluster variety coincide up to codimension $2$ with the open Calabi–Yau surfaces deformation-equivalent to the open Calabi–Yau surface $U$ associated to the asymptotic configuration of 7-branes by Construction 3.10.
@@
To prove that the map $F$ is injective, let $A_{1}$ and $A_{2}$ be two asymptotic configurations of 7-branes such that the corresponding open Calabi–Yau surfaces $U_{A_{1}}$ and $U_{A_{2}}$ are isomorphic. For every $k\in\{1,2\}$, we have $U_{A_{k}}=Y_{A_{k}}\setminus D_{A_{k}}$, where $(Y_{A_{k}},D_{A_{k}})$ is a log Calabi–Yau surface defined via a toric model $(Y_{A_{k}},D_{A_{k}})\rightarrow(\overline{Y}_{A_{k}},\overline{D}_{A_{k}})$. By [58, Proposition 3.27] (see also [24, Theorem 1]), any two toric models of a log Calabi–Yau surface are connected by a sequence of corner blow-ups/blow-downs and elementary cluster transformations. Since elementary cluster transformations correspond to Hanany–Witten moves at the level of the asymptotic configurations of 7-branes, it follows that $A_{1}$ and $A_{2}$ are equivalent in the sense of Definition 3.2. This concludes the proof that the map $F$ is injective. ∎
@@ 3.2. Webs of 5-branes with 7-branes
Figure 6. A web of 5-branes with 7-branes
@@
Let $\overline{W}\subset\mathbb{R}^{2}$ be a web of 5-branes, and $\mathbf{a}:=((a_{e,i})_{1\leq i\leq n_{e}})_{E\in L(\overline{W})}$ a 7-brane data for $\overline{W}$. We define an integral affine manifold with singularities $B$, by introducing for every leg $e$ of $\overline{W}$, $n_{e}$ focus-focus singularities $x_{e,1},\dots,x_{e,n_{e}}$ on $e$ with monodromy-invariant direction the direction $(p_{e},q_{e})$ of $e$, ordered as in Figure 6. Let $W\subset B$ be the embedded graph in $B$ obtained from $\overline{W}$ as follows:
@@
    Finally, we assign weight $w_{e}$ to the edge $v,x_{e,1}]$, weights $w_{e}-\sum_{j=1}^{i}a_{e,j}$ to the edge $x_{e,i},x_{e,i+1}]$ for all $1\leq i\leq n_{e}-1$, and weight $w_{e}-\sum_{j=1}^{n_{e}}a_{e,j}$ to the leg $x_{e,n_{e}}+\mathbb{R}_{\geq 0}(p_{e},q_{e})$ if $\sum_{i=1}^{n_{e}}a_{e,i}<w_{e}$.
@@ 3.3. Asymptotic webs of 5-branes with 7-branes and log Calabi–Yau surfaces
Recall from §2.2 that the asymptotic 5-brane $\overline{W}^{\mathrm{asym}}$ determines a toric surface $(\overline{Y},\overline{D})$ with an ample line bundle $\overline{L}$. Every leg $e\in L(\overline{W}^{\mathrm{asym}})$ corresponds to a toric irreducible component $\overline{D}_{e}$ of $\overline{D}$. Denote by $\overline{D}_{e}^{\circ}$ the open stratum of $\overline{D}_{e}$, that is, the complement in $\overline{D}_{e}$ of the 0-dimensional strata of $\overline{D}$. For every leg $e\in L(\overline{W}^{\mathrm{asym}})$, pick $n_{e}$ distinct points $y_{e,i}$ of $\overline{D}_{e}^{\circ}$, where $1\leq i\leq n_{e}$. Let $Y$ be the projective surface obtained by blowing up the toric surface $\overline{Y}$ at the points $y_{e,i}$ for all $e\in L(\overline{W}^{\mathrm{asym}})$ and $1\leq i\leq n_{e}$. Then, denoting by $D$ the strict transform of $\overline{D}$ in $Y$, with irreducible components $D_{e}$, the pair $(Y,D)$ is a log Calabi–Yau surface. Finally, we define a line bundle $L$ on the surface $Y$ by
@@
where $\pi:Y\rightarrow\overline{Y}$ is the blow-up map, and $E_{e,i}=\pi^{-1}(y_{e,i})$ are the exceptional divisors in $Y$ over the points $y_{e,i}$ – see Figure 7.
Figure 7. The log Calabi–Yau surface associated to an asymptotic web of 5-branes with 7-branes.
@@
For Type I webs $W^{\mathrm{asym}}$, we do not impose any restrictions on the configuration of the points $y_{e,i}$. In particular, by [51, Proposition 2.9], one can assume that $(Y,D)$ is generic in its deformation class, in the sense that there does not exist any effective line bundle $L^{\prime}$ on $Y$ whose restriction $L^{\prime}|_{D}$ is trivial.
For Type II webs, we require that the points $y_{e,i}$ satisfy the Menelaus condition with respect to $L$, that is, that the line bundle $L|_{D}$ is trivial – see for example [85] for a previous use of this terminology. By [51, Proposition 2.9], one can then assume that $(Y,D)$ is generic among its deformations preserving the Menelaus condition with respect to $L$, in the sense that the only effective line bundles $L^{\prime}$ on $Y$ with trivial restriction $L^{\prime}|_{D}$ are proportional to $L$.
@@
Since $D$ is a nodal curve of arithmetic genus one, the group of $\mathrm{Pic}^{0}(D)$ of line bundles on $D$ with degree 0 on every irreducible component of $D$ is isomorphic to $\mathbb{C}^{\star}$. By Equation 3.5, we have $L|_{D}\in\mathrm{Pic}^{0}(D)$, and so the Menelaus condition can be written as $L|_{D}=1\in\mathbb{C}^{\star}$. Recall that the points $y_{e,i}$ lie in the interior of the big torus orbits in the boundary components $\overline{D}_{e}$’s of $\overline{D}$. Denote by $z_{e}$ a coordinate on the component $\overline{D}_{e}$, with the convention that we choose an orientation $\overline{D}$ and denote the coordinates of the two nodal points intersecting adjacent components by 0 and $\infty$ respectively, following this orientation. Denoting by $z_{e}(y_{e,i})\in\mathbb{C}^{*}$ the coordinates of $y_{e,i}$, the Menelaus condition amounts to the requirement:
@@ 3.4. Webs of $5$-branes with $7$-branes and degenerations of log Calabi–Yau surfaces
In this section, we explain how a web of $5$-branes with $7$-branes $W$ defines a one-parameter degeneration $(\mathcal{Y},\mathcal{D})\to\mathbb{C}$ of the log Calabi–Yau surface $(Y,D)$, associated to $W^{\mathrm{asym}}$ as in §3.2. We also explain how to construct a degeneration $\mathcal{L}$ of the line bundle $L$ on $Y$.
Let $W$ be a web of $5$-branes with $7$-branes, obtained from a web of $5$ branes $\overline{W}$ which is dual to an integral convex piecewise linear function as in §2.2.3. As reviewed in §2.2.3, the web $\overline{W}$ defines a polarized toric degeneration
@@
with general fiber the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ defined by the associated asymptotic web $\overline{W}^{\mathrm{asym}}$ as in §2.2.2.
@@
where $p:\mathcal{Y}\rightarrow\overline{\mathcal{Y}}$ is the blow-up map, and $\mathcal{E}_{e,i}$ are the exceptional divisors in $\mathcal{Y}$ over the images of the sections $\sigma_{e,i}$. By construction, the restriction of $\mathcal{L}$ to a general fiber $(\mathcal{Y}_{t},\mathcal{D}_{t})$, which is deformation equivalent to $(Y,D)$, is deformation equivalent to the line bundle $L$ defined in Equation (3.3). From now on we refer to